OpenStudy (anonymous):

Solve for x: (3x-9)(x-1)=0

OpenStudy (anonymous):

3x-9=0 OR x-1=0 then?

OpenStudy (anonymous):

when u distribute, u get 3x-3x-9x+9=0, which is then -9x+9=0 change the 9 to the other side f the equals sign to make it -9x=-9 divide both side by -9, to get x=1